# Evidence-Decision-Feedback: Theory-Driven Adaptive Scaffolding for LLM Agents

### 29 Mar 2026 v3

@@ Abstract
LLMs offer tremendous opportunity for pedagogical agents to help students construct knowledge and develop problem-solving skills, yet many of these agents operate on a “one-size-fits-all” basis, limiting their ability to personalize support. To address this, we introduce Evidence-Decision-Feedback (EDF), a theoretical framework for adaptive scaffolding with LLM agents. EDF integrates elements of intelligent tutoring systems (ITS) and agentic behavior by organizing interactions around evidentiary inference, pedagogical decision-making, and adaptive feedback. We instantiate EDF through Copa, a Collaborative Peer Agent for STEM+C problem-solving. In an authentic high school classroom study, we show that EDF-guided interactions align feedback with students’ demonstrated understanding and task mastery; promote scaffold fading; and support interpretable, evidence-grounded explanations without fostering overreliance.
@@ 1 Introduction
We developed a collaborative peer agent, Copa, to instantiate EDF by mapping each EDF component to concrete elements of Copa’s multi-agent architecture (Section 4). In an authentic high school science classroom with $n=33$ dyads, students learned physics and computing by building computational models in C2STEM, an open-ended learning environment, where we deployed Copa to support exploration and inquiry during problem solving [17]. Our findings suggest that EDF supports personalized, effective, and interpretable adaptive scaffolding without fostering overreliance on the agent, advancing the development of theory-aligned, agentic LLM systems in AIED.
@@ 2 Background
Table 1 compares our proposed framework to recent alternatives across several dimensions. In contrast to intelligent tutoring systems (ITS) that emphasize direct instruction and optimized learning paths, open-ended learning environments (OELEs) emphasize exploration- and inquiry-based problem solving, in which students construct knowledge through interaction with complex scenarios (e.g., “learning-by-modeling”) rather than solely through information delivery [3].
Building effective agents for these environments necessitates robust learner modeling to support scaffolding that is both personalized to students and adaptive over time. This requires (1) attention to factors beyond cognition alone (e.g., self-regulation); (2) dialogic feedback grounded in social-cognitive and social-constructivist principles; and (3) that dialogue is treated as a first-class support mechanism, shaping what the agent says, when it probes, and how it responds. Additionally, agent decision-making and feedback must be interpretable to stakeholders and traceable through the system to its theoretical foundations.
ITS frameworks are designed primarily for well-structured domains, constrained solution paths, and tightly specified learner actions, but are less suited to OELEs where student progress is nonlinear, strategies vary, and evidence of understanding is distributed across evolving dialogue, behavior, and environment artifacts. These frameworks offer stronger theoretical grounding, interpretability, and learner modeling, but often lack dialogue-centered decision making and dialogic support for open-ended interaction. LLMs help address this gap by enabling agents to interpret richer learner evidence, engage in natural language dialogue, and generate context-sensitive feedback that adapts to students’ evolving needs. However, many existing LLM-based agents underutilize these capabilities, relying on feedback limited in personalization, adaptivity, or multi-turn interaction [19, 24, 34]. Simultaneously, advanced systems that adopt agentic or multi-agent architectures often provide less guidance on how scaffolding is grounded in pedagogy, straying from ITS theoretical foundations [22, 6, 14].
@@ 4 Copa
Copa is embedded within the C2STEM [17] learning environment (see Figure 2), which targets 1- and 2-D kinematics through open-ended dyadic computational modeling tasks. Students simulate the motion of objects such as a truck (1-D) and a drone (2-D) using kinematic equations and block-based code, interacting with Copa as they see fit. C2STEM generates rich log data to support agent reasoning, including logged actions (e.g., editing code blocks), model state (the code blocks on screen), and task context (the model component currently being addressed, e.g., variable initialization). Logged actions are further mapped to processed actions to provide the LLM with a semantically meaningful representation (e.g., set_blk1234_vel_4 $\rightarrow$ Set Velocity to 4 m/s). Log data also allows us to track students’ progress by defining rubric criteria that must be met for each task. Importantly, we distinguish between students’ understanding—their verbalized knowledge to the agent—and their task mastery—performance measured as the percentage score achieved for a given C2STEM task.
@@
All agents have specific theoretical motivations. For example, the KnowledgeAgent estimates students’ ZPD and identifies the “immediate next step” needed based on task progress, while the DialogueAgent is instructed to “promote self-efficacy through encouragement and praise” (SCT) and ask “probing questions” (Social Constructivism). All task rubrics, as well as agent prompts, input features and their values, outputs, functionality, and theoretical foundations are detailed in our Supplementary Materials [^1] https://github.com/claytoncohn/AIED26_Supplementary_Materials.
@@ 5 Evaluation
    Is students’ verbalized understanding aligned with their task mastery when interacting with Copa?
@@ 5.1 RQ1: Scaffolding Adaptivity
Table 2: Dialogue policy adaptation across task mastery quintiles.
@@ 5.4 RQ4: Feedback Interpretability
Link 2, referred to as Alignment, assesses semantic coherence between Copa’s Decision (dialogue policy selection) and extracted Evidence. We compute semantic similarity using Sentence-BERT embeddings (SBERT; all-MiniLM-L6-v2)—selected for its balance of size and performance—by encoding Copa’s dialogue policy and dialogue state summaries, then comparing them via cosine similarity. Significance is evaluated via a permutation test (n=1000n=1000) [^5] Larger $n$ due to fewer utterances relative to environment actions in Link 1., comparing mean similarity of true evidence-policy pairs to shuffled baselines. Link 3, referred to as Faithfulness, evaluates whether Copa’s final Feedback adheres to its Decision. We assess faithfulness by again computing the SBERT-based cosine similarity between the dialogue policy and talk move, using the same permutation testing procedure as Link 2.
@@ 6 Discussion and Conclusions
The primary limitations of our work include evaluating scaffold adaptation solely as a function of task mastery rather than a broader set of learner attributes, and the absence of an RCT due to institutional constraints: reported relationships are correlational rather than causal. Our evaluation focuses on within-system behavior rather than comparison against alternatives, and although EDF is a general framework, we evaluate it through only one implementation, Copa. Researcher-led classroom instruction may introduce effects that would not arise under independent teacher implementation. Manual triangulation across log, screen, and video data suggests Copa supports self-regulation and produces pedagogically sound reasoning chains; however, rigorous stakeholder-centric evaluation across these axes remains future work.
@@ 6.0.1 Acknowledgements
We thank the Institute of Education Sciences (IES; award R305C240010) and National Science Foundation (NSF; awards IIS-2327708 and DRL-2112635) for their support. The opinions expressed are those of the authors and do not represent the views of the IES or NSF.

### 31 Mar 2026 v4

@@ Abstract
LLMs offer tremendous opportunities for pedagogical agents to help students construct knowledge and develop problem-solving skills, yet many of these agents operate on a “one-size-fits-all” basis, limiting their ability to personalize support. To address this, we introduce Evidence-Decision-Feedback (EDF), a theoretical framework for adaptive scaffolding with LLM agents. EDF integrates elements of intelligent tutoring systems (ITS) and agentic behavior by organizing interactions around evidentiary inference, pedagogical decision-making, and adaptive feedback. We instantiate EDF through Copa, a Collaborative Peer Agent for STEM+C problem-solving. In an authentic high school classroom study, we show that EDF-guided interactions align feedback with students’ demonstrated understanding and task mastery; promote scaffold fading; and support interpretable, evidence-grounded explanations without fostering overreliance.
@@ 1 Introduction
We developed a collaborative peer agent, Copa, to instantiate EDF by mapping each EDF component to concrete elements of Copa’s multi-agent architecture (Section 4). In an authentic high school science classroom with $n=33$ dyads, we deployed Copa to support exploration and inquiry during problem solving within an open-ended learning environment (OELE), C2STEM [17]. Our findings suggest that EDF supports personalized, effective, and interpretable adaptive scaffolding without fostering overreliance on the agent.
@@ 2 Background
In contrast to intelligent tutoring systems (ITS) that emphasize direct instruction and optimized learning paths, OELEs emphasize exploration- and inquiry-based problem solving, in which students construct knowledge through interaction with complex scenarios (e.g., “learning-by-modeling”)  [3]. Building effective agents for these environments necessitates robust learner modeling to support scaffolding that is both personalized to students and adaptive over time. This requires (1) attention to factors beyond cognition alone (e.g., self-regulation); (2) dialogic feedback grounded in social-cognitive and social-constructivist principles; and (3) that dialogue is treated as a first-class support mechanism, shaping what the agent says, when it probes, and how it responds. Additionally, agent decision-making and feedback must be interpretable to stakeholders and traceable to its theoretical foundations.
Table 1 compares our proposed framework to recent alternatives across several dimensions. ITS frameworks are designed primarily for well-structured domains, constrained solution paths, and tightly specified learner actions, but are less suited to OELEs where student progress is nonlinear, strategies vary, and evidence of understanding is distributed across evolving dialogue, behavior, and environment artifacts. These frameworks offer stronger theoretical grounding, interpretability, and learner modeling, but often lack dialogue-centered decision making and dialogic support for open-ended interaction.
LLMs help address this gap by enabling agents to interpret richer learner evidence, engage in natural language dialogue, and generate context-sensitive feedback that adapts to students’ evolving needs. However, many existing LLM-based agents underutilize these capabilities, relying on feedback limited in personalization, adaptivity, or multi-turn interaction [19, 24, 34]. Simultaneously, advanced systems that adopt agentic or multi-agent architectures often provide less guidance on how scaffolding is grounded in pedagogy [22, 6, 14].
@@ 4 Copa
Copa is embedded within the C2STEM learning environment (see Figure 2), which targets 1- and 2-D kinematics through open-ended dyadic computational modeling tasks. Students simulate the motion of objects such as a truck (1-D) and a drone (2-D) using kinematic equations and block-based code, interacting with Copa as they see fit. C2STEM generates rich log data to support agent reasoning, including logged actions (e.g., editing code blocks), model state (the code blocks on screen), and task context (the model component currently being addressed, e.g., variable initialization).
Logged actions are mapped to processed actions to provide the LLM with a semantically meaningful representation (e.g., set_blk1234_vel_4 $\rightarrow$ Set Velocity to 4 m/s). Logs also allow us to track student progress by defining rubric criteria for each task. Importantly, we distinguish between students’ understanding—their verbalized knowledge to the agent—and their task mastery—performance measured as the percentage score achieved for a given task.
@@
All agents have specific theoretical motivations. For example, the KnowledgeAgent estimates students’ ZPD and identifies the “immediate next step” needed based on task progress, while the DialogueAgent is instructed to “promote self-efficacy through encouragement and praise” (SCT) and ask “probing questions” (Social Constructivism). All task rubrics, as well as agent prompts, input features and their values, outputs, functionality, and theoretical foundations are detailed in our Supplementary Materials [^1] https://github.com/claytoncohn/AIED26˙Supplementary˙Materials.
@@ 5 Evaluation
    Does students’ verbalized understanding align with their task mastery when interacting with Copa?
@@ 5.1 RQ1: Scaffolding Adaptivity
Table 2: Dialogue policy adaptation across task mastery quintiles.
@@ 5.4 RQ4: Feedback Interpretability
Link 2, referred to as Alignment, assesses semantic coherence between Copa’s Decision (dialogue policy selection) and extracted Evidence. We compute semantic similarity using Sentence-BERT embeddings (all-MiniLM-L6-v2)—selected for its balance of size and performance—by encoding Copa’s dialogue policy and dialogue state summaries, then comparing them via cosine similarity. Significance was evaluated via a permutation test (n=1000n=1000) [^5] Larger $n$ due to fewer utterances relative to environment actions in Link 1., comparing mean similarity of true evidence-policy pairs to shuffled baselines. Link 3, referred to as Faithfulness, evaluates whether Copa’s final Feedback adheres to its Decision. We assessed faithfulness by computing the SBERT-based cosine similarity between the dialogue policy and talk move, using the same Link 2 testing procedure.
Table 3: Interpretability results for all three EDF links.
@@ 6 Discussion and Conclusions
The primary limitations of our work include evaluating scaffold adaptation solely as a function of task mastery rather than a broader set of learner attributes, and the absence of an RCT due to institutional constraints: reported relationships are correlational rather than causal. Our evaluation focuses on within-system behavior rather than comparison against alternatives, and although EDF is a general framework, we evaluate it through Copa only. Manual triangulation across log, screen, and video data suggests Copa supports self-regulation and produces pedagogically sound reasoning chains; however, rigorous stakeholder-centric evaluation across these axes remains future work.
@@ 6.0.1 Acknowledgements
We thank the Institute of Education Sciences (IES; award R305C240010) and National Science Foundation (NSF; awards IIS-2327708 and DRL-2112635) for their support. All opinions are those of the authors, not the IES or NSF.
[…]
The authors have no competing interests to declare.